The partition coefficient of a base B (monoacidic) at a particular pH, in an octanol-water mixture
is 1.45. Consider an experiment where 3.5 g of B are shaken with 100 mL of octanol and 50 mL
of water.
Molar mass of B = 101.19 g mol-1
(i) Calculate the mass of B in the aqueous layer after shaking. (4 marks)
(ii) Calculate the volume of 0.5 M HCl required to neutralize the aqueous layer
1
Expert's answer
2020-06-08T15:25:12-0400

Partition Coefficient K = [concentration in octanol] / [concentration in water]

Now

1.45 = [mass in octanol / 100]/ [mass in water/ 50]

mass in octanol / mass in water = 2.9

Mass in octanol = 2.9 × mass in water

Total mass = 3.5 g

3.9× mass in water = 3.5 g

Mass in water = 3.5/ 3.9

= 0.9 g

Molarity of the aqueous solution

= (0.9/ 101.19)/ 0.05

= 0.177 M

So this is the base concentration

Here 50 ml of 0.177M of water solution present so to neutralize volume of HCl required

= (0.177 × 50) / 0.5

=17.7 ml